OpenStudy (anonymous):

Which is a factor of 2x2+ 8x − 90? x + 2 x − 15 x + 9 x − 4

OpenStudy (lochana):

mm.. may be x+9

OpenStudy (anonymous):

\(\large 2x^2+8x-90\)\(\large =2(x^2+4x-45)\\\large =2(x^2+9x-5x-45)\\\large =2[x(x+9)-5(x+9)]\\\large =2(x-5)(x+9)\) Thus, \(\large x + 9\) is a factor